# Critical behavior and multistability in quasi-two-dimensional turbulence

### v1

@@ Appendix D Calibration of force and velocity
The voltage between the device electrodes and the resulting current differ from the voltage applied to the primary winding of the center-tapped transformer and the current measured by the lock-in amplifier due to circuit losses. These losses need to be taken into account to obtain accurate velocities and driving pressure gradients. The frequency and applied voltage dependence of the circuit losses were characterized by directly measuring the oscillating voltage across the device with a separate lock-in amplifier. The current losses were characterized by injecting a known oscillating current immediately after the device on the bias side (see Fig. S1(c)). The resulting transfer functions were used to correct the driving force (proportional to the voltage across the device) and the flow velocity (proportional to the detected current).
@@
The calculation of the mechanical quantities also requires the effective spring constant of the deformation of the substrate in the basin region. Let $C_{0}$ be the capacitance of the device corresponding to the zero DC bias voltage $U_{B}$. We characterize the effective spring constant $k_{p}$ by measuring the dependence of the device capacitance on the applied bias voltage $U_{B}$ similarly to ref. Souris et al. (2017). The force acting on the basin plates is
$F=\dfrac{CU^{2}}{2D}=k_{p}x,$ (S18)
@@
$C=\dfrac{A\varepsilon_{0}}{D-2x}\approx\dfrac{A\varepsilon_{0}}{D}+\dfrac{2\varepsilon_{0}A}{D^{2}}x=C_{0}+\dfrac{2C_{0}}{D}x,$ (S19)
where $A$ is the basin area, $\varepsilon$ is the vacuum permittivity, and we assumed $x\ll D$, taking into account only the first two terms of Taylor expansion. Furthermore, $x$ can be expressed from Eq. (S18) and substituting into Eq. (S19) we get
$C=C_{0}+\sigma U^{2},$ (S20)
@@
$C_{corr}=1/(1/C_{meas}-1/100\textrm{nF})-2C_{\mathtt{channel}}.$ (S21)
@@
The corrected $C_{corr}$ as a function of the applied voltage squared $U^{2}$ is shown in Fig. S5. The data fit with the Eq. (S20) to obtain $C_{0}$ and $k_{p}$, shown in Tab. 2
@@
Fig.  S5: The corrected capacitance of all four Helmholtz resonators against the applied voltage. Only the DC bias voltage amplitude was changed, and the amplitude of the AC drive was kept constant at 0.02 Vrms. The black dashed line is a fit of the Eq. (S20). From the fit parameters we calculate $C_{0}$ and $k_{p}$ and these values can be found in the Tab. 2.
@@
Corner 1.24 338.6
Rounded corner 1.19 312.9
Jet 1.45 335.7
Grid 1.62 466.4
[…]
The relevant mechanical parameters are the pressure gradient $\delta P/l$ along the flow channel resulting from the electrostatic driving force and the amplitude of the flow velocity. The equation for $\delta P/l$ can be obtained from the driving term in (S11)
$\dfrac{\delta P}{l}=\dfrac{2AC_{0}U_{B}}{l(\chi V_{B}k_{p}+2A^{2})h}U_{0},$ (S22)
where $U_{0}(t)$ was replaced by its amplitude $U_{0}$. Based on ref. Varga et al. (2020), velocity is calculated as
$v_{\mathrm{s}}=\dfrac{I}{gU_{B}C_{0}},$ (S23)
$g=\dfrac{2a\rho_{s}}{V_{B}\rho}\left(1-\dfrac{2(\varepsilon-1)}{\varepsilon}\right)\dfrac{1}{1+2\Sigma},$ (S24)
where $\varepsilon$ is the dielectric constant taken from Brooks and Donnelly (1977), and $\Sigma=\chi Dk_{p}/(2A)$.
Appendix E Fourth sound damping
Following the analysis of refs. Backhaus and Backhaus (1997); Souris et al. (2017), linear damping of the Helmholtz resonance (S8) in the temperature range 1.3 K to 2 K has two dominant contributions: viscous normal fluid motion and thermoviscous dissipation. For the inverse of resonance width $\Delta_{n}$ caused by the viscous normal flow we can write Souris et al. (2017)
$\Delta_{n}^{-1}=2\pi\dfrac{8\eta}{(D/2)^{2}}\dfrac{\rho_{s}}{\rho_{n}^{2}}\dfrac{1}{\omega_{0}^{2}},$ (S25)
where $\eta$ is the dynamic viscosity.
The second contribution is due to the heat transfer (resulting in energy loss of the oscillating flow) from the basin to the surroundings (for further details see ref. Backhaus and Backhaus (1997)). The temperature gradient between the basin and the bulk 4He is a consequence of a mechanocalorical effect, that is, the change in the superfluid density inside the basin results in a change in temperature. This effect can be characterized by the parameter
$\Gamma^{2}=\dfrac{\rho^{2}s^{2}TA^{2}(1+\Sigma)}{C_{\mathrm{th}}k_{p}},$ (S26)
where $C_{\mathrm{th}}=C_{s}V_{B}$ is the total heat capacity of 4He inside the basin ($C_{s}$ is the specific heat capacity of 4He) and $s$ is the specific entropy. For the case $\Gamma^{2}\ll 1$ (in our case $\Gamma\approx 0.22$ at 1.65 K), the influence of the additional pressure gradient caused by the presence of the temperature gradient (i.e., the fountain pressure) is negligible and for damping due to the heat flux out of the basin, it can be shown Souris et al. (2017)
$\Delta_{th}^{-1}=\dfrac{1+\Phi^{2}}{\Phi\Gamma^{2}}\left(1+\dfrac{\Gamma^{2}}{2(1+\Phi^{2})}\right)\dfrac{2\pi}{\omega_{0}},$ (S27)
where $\Phi=1/(\tau_{\mathrm{th}}\omega_{0})$. Relaxation time $\tau_{\mathrm{th}}=C_{\mathrm{th}}R_{\mathrm{th}}$ represents how quickly the system returns to thermal equilibrium. If $1/\tau_{th}<\omega_{0}$, the energy loss due to heat transport must be taken into account. The basin has two channels to exchange heat with the surroundings: helium flow through the channels or thermal conduction through the substrate. The thermal conductivity through helium in the flow channels is negligible Souris et al. (2017). The dominant limiting factor for heat conduction through the substrate is the interfacial thermal resistance (Kapitza resistance) $R_{k}$. The temperature dependence for the interface between liquid helium and aluminum is $R_{k}=28.57T^{-4.21}/A$ Souris et al. (2017), where $A$ is the basin area in cm2. Taking into account both basin plates, the total thermal resistance is $R_{\mathrm{th}}=R_{k}/2$. It should be noted, however, that the Kapitza resistance is sensitive to surface preparation, and considerable scatter exists in the reported empirical values.
The total resonance full width at half maximum (FWHM) is given by the sum of (S26) and (S27). Due to the end effects of the flow in the channels (i.e., the finite flow in the basin), the temperature oscillation inside the basin will not be uniform. Therefore, we use an effective basin area such that $A\to A/\lambda$, where $\lambda$ is an adjustable parameter. The only adjustable parameters are the scaling parameter $\lambda$ and the temperature correction (the same as for the frequency, i.e., parameter $b$ in Tab. 1). Best-fit values are listed in Tab. 1 for the four geometries.
The resulting fit to the widths of the experimentally measured resonance curves is essentially independent of the geometry of the channel, and thus Fig. S3(b) shows the average of the four fits as a single dashed line. The discrepancy between measured data and theoretical prediction is mainly in the low-temperature range; the theory underestimates the measured value by approximately a factor of two. This is likely due to uncertainty in Kapitza resistance (which is also reflected in the relatively large value of $\lambda$ in Tab. 1) since it strongly depends on the purity of the Al electrode and the structure of its surface.
Appendix F Numerical simulations
In order to better understand the behavior of the confined oscillatory flow, we performed two complementary sets of independent numerical simulations. First, a classical fluid dynamics calculation of confined incompressible turbulence based on the $k-\varepsilon$ model Davidson (2015); and second, a vortex filament simulation Hanninen and Baggaley (2014) of individual quantized vortices.
F.1 $k-\varepsilon$ simulation
The $k-\varepsilon$ is a simplified turbulence model based on the eddy viscosity approach Davidson (2015). It is a classical continuous model that does not contain any features of superfluid dynamics (e.g., mutual friction or the discreteness of superfluid vorticity), and the obtained results, therefore, have only an illustrative character. Turbulent flows of superfluid helium with large vortex densities polarized in large structures are, however, known to behave quasi-classically Babuin et al. (2014).
[…]
@@ F.1 $k-\varepsilon$ simulation
To obtain time-averaged behavior, 20 periods of the oscillatory flow were simulated with velocity amplitude $v_{0}=8\times 10^{-4}$ m/s and the frequency $f=\omega/(2\pi)=2000$ Hz, for the four geometries, which correspond to the typical frequency of the Helmholtz mode and velocity in the turbulent regime. Since the experiment measures behavior averaged over many oscillations, the calculated vorticity was averaged over the last period of the calculation (when the flow was already in an oscillatory steady state). The results are shown in Fig. S6. As can be seen in the cases of C, J, and G, large vortical structures are observed, which are absent in the case of R. These structures most likely correspond to the acoustic streaming patterns Duda et al. (2017). Note that the $k-\varepsilon$ model cannot simulate the inverse cascade since it assumes isotropic dissipation Davidson (2015). However, the results show that over long time periods, the acoustic flow drives a particular mode of the spatial vorticity distribution, which is consistent with the dynamical model presented in the main text.
@@
Fig.  S6: Results of quasi 2D finite element numerical simulation in the framework of $k-\varepsilon$ turbulence model for all four chip geometries used in the experiment. We display the vorticity of fluid averaged over one period of oscillation.
F.2 Vortex filament model simulations
@@
The motion of individual quantized vortices in uniformly oscillating superflow and stationary normal fluid confined between two walls was numerically simulated using the vortex filament model Hanninen and Baggaley (2014). Quantized vortices are represented as infinitesimally thin spatial curves, which move and deform under externally applied potential superflow and the superfluid velocity field $\mathbf{v}_{s}(\mathbf{r})$ induced by the vortices given by the Biot-Savart integral,
[…]
@@
$\frac{\partial{\mathbf{s}}}{\partial{t}}=\mathbf{v}^{\prime}_{s}+\alpha\mathbf{s}^{\prime}\times(\mathbf{v}^{\prime}_{n}-\mathbf{v}_{s})-\alpha^{\prime}\mathbf{s}^{\prime}\times\left[\mathbf{s}^{\prime}\times(\mathbf{v}^{\prime}_{n}-\mathbf{v}_{s})\right],$ (S29)
where $\alpha$ and $\alpha^{\prime}$ are temperature-dependent mutual friction coefficients Donnelly and Barenghi (1998) and $\mathbf{v}^{\prime}_{s}$ is the standard de-singularization of the Biot-Savart integral (S28) for $\mathbf{r}$ on a vortex line Hanninen and Baggaley (2014)
[…]
@@
The vortices are discretized to a set of connected points whose spacing is dynamically maintained between $10^{-8}$ and $1.2\times 10^{-8}$ m (simulations were checked with half the discretization length and yielded identical results). The time evolution was implemented using the 4th order Runge-Kutta stepping with fixed time step $\delta t=10^{-10}$ s. Planar walls placed at $z=0$ and $z=D$ are considered to be uniformly pinning; a discretization point of a vortex attached to the wall is not allowed to move. However, the vortex can reconnect with its image on the other side of the wall, which exists because of boundary conditions of no flow normal to the wall. Depinning is handled entirely by modeling pinning and reconnections, ensuring self-consistency with the rest of the model. As a result, for a sufficiently sharp angle at the contact between the vortex and the wall, the vortex can effectively slide in a manner similar to ref Nakagawa et al. (2023).
@@
We studied the evolution of two particular scenarios with mutual friction constants corresponding to 1.45 K, $\alpha=0.061$ and $\alpha^{\prime}=1.746\times 10^{-2}$ and $D=500$ nm – a half-loop attached to the bottom wall and a pair of reconnecting antiparallel vortex lines spanning the two walls. For all simulated cases, the imposed velocity was set to $v_{s}=0.25$ m/s along the x-axis and $v_{n}=0$. The time evolutions of the vortex configurations are shown in Fig. S7 for a pair of antiparallel vortices and Figs. S8, S9 for half-loops attached to the bottom wall oriented parallel or antiparallel with the imposed superfluid velocity $\mathbf{v}_{s}$. All initial vortex configurations studied annihilate in a few tens of microseconds with a typical flow oscillation period of 0.5 ms (see Fig. S3(a)). These results suggest that the vortex configuration will remain quasi-two-dimensional, with only transient three-dimensional structures short-lived compared to the flow oscillation period.
@@
Fig.  S7: Typical evolution of two colliding vortices spanning the 500 nm gap between the top and bottom walls. As an initial condition, two straight vortices of opposite signs separated by 500 nm along the $y$ direction were used. The vortices are subject to a steady flow of the superfluid component with a velocity of 25 cm/s along the $x$ axis. The figure shows several snapshots leading up to the annihilation (time given by the color), with 25 ns between subsequent snapshots. The entire evolution from the initial condition to the final annihilation takes about 40 $\mu$s (the entire lead-up to the reconnection is not shown).
@@
Fig.  S8: Annihilation of vortex half-ring with 500 nm diameter attached to the bottom wall with orientation parallel with the imposed superflow oriented along the positive x-axis direction. The initial condition was a half-circle lying in the $x=0$ plane.
@@
Fig.  S9: As in Fig. S8 with the ring orientation anti-parallel with the imposed $\mathbf{v}_{s}$. The vortex propagates in the negative $x$ direction because its self-induced velocity exceeds the imposed superfluid velocity.
Appendix G Supplementary data
In the following, we provide an overview of the complete dataset of the measured flow velocity as a function of the driving pressure gradient (Fig. S10) or, equivalently, the vortex density as a function of the flow velocity (Figs. S11, S12, S14, and S13). For device R, Fig. S14, intermittent DC resistance in the range of several MOhm was observed across the resonator electrodes, which shunts the capacitance of the device, resulting in an unreliable calibration of force and velocity. Therefore, the comparison with the other three devices ought to be limited to qualitative features of the force-velocity, or equivalently, vortex density-velocity curves.
Figure S15 shows the decreasing lifetime of the upper state in the intermittent region of device J (complementary data to Fig. 4 in the main text), which similarly shows a power law behavior, although with a steeper critical exponent of approximately 3. Finally, Fig. S16 shows the histogram of the lower state lifetimes for three temperatures and a range of drive amplitudes corresponding to mean lifetimes of 0.1 to 0.2 s. The distribution is exponential, indicating that the switching process is memoryless. Finally, Fig. S17 shows an example of the time dependence in the intermittent switching region of device J at temperature 1.65 K.
Fig.  S10: Dependence of the superfluid velocity on the pressure gradient for all four used resonators. Both quantities are calculated from measured current and applied voltage using Eqs. (S23) and (S22). Data from different temperatures (indicated by the color) are offset for clarity by a multiplicative factor of 1.1. A, corner device; B, jet device; C, rounded corner device and D, grid device.
Fig.  S11: Vortex line density as a function of velocity for all studied temperatures (indicated in the panels) for the device C. Lines connect points from a single continuous ramp of the drive. Blue – drive is ramped up, orange – drive is ramped down.
Fig.  S12: As Fig. S11, device J.
Fig.  S13: As Fig. S11, device G.
Fig.  S14: As Fig. S11, device R. Note that the electrodes of this device were intermittently shunted by DC resistance, thus making the calibration of force and velocity unreliable.
Fig.  S15: Lifetimes of the upper state in Fig. 4 of the main text as a function of flow velocity amplitude shifted by the critical velocity $v_{c}$. The inset shows the same data in logarithmic axes against reduced velocity $(v-v_{c})/v_{c}$. The dashed line indicates a power law with exponent -3, different from the -1.33 shown by the lower state lifetimes in Fig. 4 of the main text.
Fig.  S16: Histrogram of the turbulent state lifetimes in the transitional region of device J at the three indicated temperatures for datasets with a mean lifetime in the range 0.1 to 0.2 s. A straight line in semilogarithmic axes indicates exponentially decaying dependence.
Fig.  S17: Time dependence of the observed velocity amplitude in device J in the intermittently switching region at 1.65 K (main text Fig. 4(a), middle panel), drive pressure gradient approximately 9.5 kPa/m. Maximum and minimum velocity were normalized to 0 and 1, respectively; the orange line shows the binarized signal used for the detection of transitions.

### v2

@@ Appendix B Calibration of force and velocity
The voltage between the device electrodes and the resulting current differ from the voltage applied to the primary winding of the center-tapped transformer and the current measured by the lock-in amplifier due to circuit losses. These losses need to be taken into account to obtain accurate velocities and driving pressure gradients. The frequency and applied voltage dependence of the circuit losses were characterized by directly measuring the oscillating voltage across the device with a separate lock-in amplifier. The current losses were characterized by injecting a known oscillating current immediately after the device on the bias side (see Fig. 7(b)). The resulting transfer functions were used to correct the driving force (proportional to the voltage across the device) and the flow velocity (proportional to the detected current).
@@
The calculation of the mechanical quantities also requires the effective spring constant of the deformation of the substrate in the basin region. Let $C_{0}$ be the capacitance of the device corresponding to the zero DC bias voltage $U_{B}$. We characterize the effective spring constant $k_{p}$ by measuring the dependence of the device capacitance on the applied bias voltage $U_{B}$ similarly to ref. [24]. The force acting on the basin plates is
$F=\dfrac{CU^{2}}{2D}=k_{p}x,$ (20)
@@
$C=\dfrac{A\varepsilon_{0}}{D-2x}\approx\dfrac{A\varepsilon_{0}}{D}+\dfrac{2\varepsilon_{0}A}{D^{2}}x=C_{0}+\dfrac{2C_{0}}{D}x,$ (21)
where $A$ is the basin area, $\varepsilon$ is the vacuum permittivity, and we assumed $x\ll D$, taking into account only the first two terms of Taylor expansion. Furthermore, $x$ can be expressed from Eq. (20) and substituting into Eq. (21) we get
$C=C_{0}+\sigma U^{2},$ (22)
@@
$C_{corr}=1/(1/C_{meas}-1/100\textrm{nF})-2C_{\mathtt{channel}}.$ (23)
@@
The corrected $C_{corr}$ as a function of the applied voltage squared $U^{2}$ is shown in Fig. 9. The data fit with the Eq. (22) to obtain $C_{0}$ and $k_{p}$, shown in Tab. 1
@@
Figure 9: The corrected capacitance of all four Helmholtz resonators against the applied voltage. Only the DC bias voltage amplitude was changed, and the amplitude of the AC drive was kept constant at 0.02 Vrms. The black dashed line is a fit of the Eq. (22). From the fit parameters we calculate $C_{0}$ and $k_{p}$ and these values can be found in the Tab. 1.
The relevant mechanical parameters are the pressure gradient $\delta P/l$ along the flow channel resulting from the electrostatic driving force and the amplitude of the flow velocity. The equation for $\delta P/l$ can be obtained from the driving term in (10)
$\dfrac{\delta P}{l}=\dfrac{2AC_{0}U_{B}}{l(\chi V_{B}k_{p}+2A^{2})h}U_{0},$ (24)
where $U_{0}(t)$ was replaced by its amplitude $U_{0}$. Based on ref. [17], velocity is calculated as
$v_{\mathrm{s}}=\dfrac{I}{gU_{B}C_{0}},$ (25)
$g=\dfrac{2a\rho_{s}}{V_{B}\rho}\left(1-\dfrac{2(\varepsilon-1)}{\varepsilon}\right)\dfrac{1}{1+2\Sigma},$ (26)
where $\varepsilon$ is the dielectric constant taken from [63], and $\Sigma=\chi Dk_{p}/(2A)$.
Appendix C Fourth sound damping
Following the analysis of refs. [64, 24], linear damping of the Helmholtz resonance (7) in the temperature range 1.3 K to 2 K has two dominant contributions: viscous normal fluid motion and thermoviscous dissipation. For the inverse of resonance width $\Delta_{n}$ caused by the viscous normal flow we can write [24]
$\Delta_{n}^{-1}=2\pi\dfrac{8\eta}{(D/2)^{2}}\dfrac{\rho_{s}}{\rho_{n}^{2}}\dfrac{1}{\omega_{0}^{2}},$ (27)
where $\eta$ is the dynamic viscosity.
The second contribution is due to the heat transfer (resulting in energy loss of the oscillating flow) from the basin to the surroundings (for further details see ref. [64]). The temperature gradient between the basin and the bulk 4He is a consequence of a mechanocalorical effect, that is, the change in the superfluid density inside the basin results in a change in temperature. This effect can be characterized by the parameter
$\Gamma^{2}=\dfrac{\rho^{2}s^{2}TA^{2}(1+\Sigma)}{C_{\mathrm{th}}k_{p}},$ (28)
where $C_{\mathrm{th}}=C_{s}V_{B}$ is the total heat capacity of 4He inside the basin ($C_{s}$ is the specific heat capacity of 4He) and $s$ is the specific entropy. For the case $\Gamma^{2}\ll 1$ (in our case $\Gamma\approx 0.22$ at 1.65 K), the influence of the additional pressure gradient caused by the presence of the temperature gradient (i.e., the fountain pressure) is negligible and for damping due to the heat flux out of the basin, it can be shown that [24]
$\Delta_{th}^{-1}=\dfrac{1+\Phi^{2}}{\Phi\Gamma^{2}}\left(1+\dfrac{\Gamma^{2}}{2(1+\Phi^{2})}\right)\dfrac{2\pi}{\omega_{0}},$ (29)
where $\Phi=1/(\tau_{\mathrm{th}}\omega_{0})$. Relaxation time $\tau_{\mathrm{th}}=C_{\mathrm{th}}R_{\mathrm{th}}$ represents how quickly the system returns to thermal equilibrium. If $1/\tau_{th}<\omega_{0}$, the energy loss due to heat transport must be taken into account. The basin has two channels to exchange heat with the surroundings: helium flow through the channels or thermal conduction through the substrate. The thermal conductivity through helium in the flow channels is negligible [24]. The dominant limiting factor for heat conduction through the substrate is the interfacial thermal resistance (Kapitza resistance) $R_{k}$. The temperature dependence for the interface between liquid helium and aluminum is $R_{k}=28.57T^{-4.21}/A$ [24], where $A$ is the basin area in cm2. Taking into account both basin plates, the total thermal resistance is $R_{\mathrm{th}}=R_{k}/2$. It should be noted, however, that the Kapitza resistance is sensitive to surface preparation, and considerable scatter exists in the reported empirical values.
The total resonance full width at half maximum (FWHM) is given by the sum of (28) and (29). Due to the end effects of the flow in the channels (i.e., the finite flow in the basin), the temperature oscillation inside the basin will not be uniform. Therefore, we use an effective basin area such that $A\to A/\lambda$, where $\lambda$ is an adjustable parameter. The only adjustable parameters are the scaling parameter $\lambda$ and the temperature correction (the same as for the frequency, i.e., parameter $b$ in Tab. 1). Best-fit values are listed in Tab. 1 for the four geometries.
The resulting fit to the widths of the experimentally measured resonance curves is essentially independent of the geometry of the channel, and thus Fig. 8(b) shows the average of the four fits as a single dashed line. The discrepancy between measured data and theoretical prediction is mainly in the low-temperature range; the theory underestimates the measured value by approximately a factor of two. This is likely due to uncertainty in Kapitza resistance (which is also reflected in the relatively large value of $\lambda$ in Tab. 1) since it strongly depends on the purity of the Al electrode and the structure of its surface.
Appendix D Numerical simulations
In order to better understand the behavior of the confined oscillatory flow, we performed two complementary sets of independent numerical simulations. First, a classical fluid dynamics calculation of confined incompressible turbulence based on the $k-\varepsilon$ model [4]; and second, a vortex filament simulation [65] of individual quantized vortices.
D.1 $k-\varepsilon$ simulation
The $k-\varepsilon$ is a simplified turbulence model based on the eddy viscosity approach [4]. It is a classical continuous model that does not contain any features of superfluid dynamics (e.g., mutual friction or the discreteness of superfluid vorticity), and the obtained results, therefore, have only an illustrative character. Turbulent flows of superfluid helium with large vortex densities polarized in large structures are, however, known to behave quasi-classically [66].
@@ D.1 $k-\varepsilon$ simulation
To obtain time-averaged behavior, 20 periods of the oscillatory flow were simulated with velocity amplitude $v_{0}=8\times 10^{-4}$ m/s and the frequency $f=\omega/(2\pi)=2000$ Hz, for the four geometries, which correspond to the typical frequency of the Helmholtz mode and velocity in the turbulent regime. Since the experiment measures behavior averaged over many oscillations, the calculated vorticity was averaged over the last period of the calculation (when the flow was already in an oscillatory steady state). The results are shown in Fig. 10. As can be seen in the cases of C, J, and G, large vortical structures are observed, which are absent in the case of R. These structures most likely correspond to the acoustic streaming patterns [67]. Note that the $k-\varepsilon$ model cannot simulate the inverse cascade since it assumes isotropic dissipation [4]. However, the results show that over long time periods, the acoustic flow drives a particular mode of the spatial vorticity distribution, which is consistent with the dynamical model presented in the main text.
@@
Figure 10: Vorticity of the flow averaged over one period of oscillation in quasi 2D finite element numerical simulation based on the $k-\varepsilon$ turbulence model for all four chip geometries used in the experiment. The simulation was performed using the COMSOL FEM suite.
D.2 Vortex filament model simulations
@@
The motion of individual quantized vortices in uniformly oscillating superflow and stationary normal fluid confined between two walls was numerically simulated using the vortex filament model [65]. Quantized vortices are represented as infinitesimally thin spatial curves, which move and deform under externally applied potential superflow and the superfluid velocity field $\mathbf{v}_{s}(\mathbf{r})$ induced by the vortices given by the Biot-Savart integral,
$\mathbf{v}_{s}(\mathbf{r})=\frac{\kappa}{4\pi}\int_{L}\frac{\mathbf{s}^{\prime}\mathbf{\times}(\mathbf{r}-\mathbf{s})}{|\mathbf{r}-\mathbf{s}|^{3}}\mathrm{d}\xi,$ (30)
@@
$\frac{\partial{\mathbf{s}}}{\partial{t}}=\mathbf{v}^{\prime}_{s}+\alpha\mathbf{s}^{\prime}\times(\mathbf{v}^{\prime}_{n}-\mathbf{v}_{s})-\alpha^{\prime}\mathbf{s}^{\prime}\times\left[\mathbf{s}^{\prime}\times(\mathbf{v}^{\prime}_{n}-\mathbf{v}_{s})\right],$ (31)
where $\alpha$ and $\alpha^{\prime}$ are temperature-dependent mutual friction coefficients [25] and $\mathbf{v}^{\prime}_{s}$ is the standard de-singularization of the Biot-Savart integral (30) for $\mathbf{r}$ on a vortex line [65]
$\mathbf{v}^{\prime}_{s}(\mathbf{r})=\frac{\kappa}{4\pi}\mathbf{s}^{\prime}\times\mathbf{s}^{\prime\prime}\ln\left(\frac{2\sqrt{l_{+}l_{-}}}{e^{1/2}a}\right)+\frac{\kappa}{4\pi}\int_{L^{\prime}}\frac{\mathbf{s}^{\prime}\mathbf{\times}(\mathbf{r}-\mathbf{s})}{|\mathbf{r}-\mathbf{s}|^{3}}\mathrm{dd}\xi,$ (32)
@@
The vortices are discretized to a set of connected points whose spacing is dynamically maintained between $10^{-8}$ and $1.2\times 10^{-8}$ m (simulations were checked with half the discretization length and yielded identical results). The time evolution was implemented using the 4th order Runge-Kutta stepping with fixed time step $\delta t=10^{-10}$ s. Planar walls placed at $z=0$ and $z=D$ are considered to be uniformly pinning; a discretization point of a vortex attached to the wall is not allowed to move. However, the vortex can reconnect with its image on the other side of the wall, which exists because of boundary conditions of no flow normal to the wall. Depinning is handled entirely by modeling pinning and reconnections, ensuring self-consistency with the rest of the model. As a result, for a sufficiently sharp angle at the contact between the vortex and the wall, the vortex can effectively slide in a manner similar to ref [68].
@@
We studied the evolution of two particular scenarios with mutual friction constants corresponding to 1.45 K, $\alpha=0.061$ and $\alpha^{\prime}=1.746\times 10^{-2}$ and $D=500$ nm – a half-loop attached to the bottom wall and a pair of reconnecting antiparallel vortex lines spanning the two walls. For all simulated cases, the imposed velocity was set to $v_{s}=0.25$ m/s along the x-axis and $v_{n}=0$. The time evolutions of the vortex configurations are shown in Fig. 11 for a pair of antiparallel vortices and Figs. 12, 13 for half-loops attached to the bottom wall oriented parallel or antiparallel with the imposed superfluid velocity $\mathbf{v}_{s}$. All initial vortex configurations studied annihilate in a few tens of microseconds with a typical flow oscillation period of 0.5 ms (see Fig. 8(a)). These results suggest that the vortex configuration will remain quasi-two-dimensional, with only transient three-dimensional structures short-lived compared to the flow oscillation period.
@@
Figure 11: Typical evolution of two colliding vortices spanning the 500 nm gap between the top and bottom walls. As an initial condition, two straight vortices of opposite signs separated by 500 nm along the $y$ direction were used. The vortices are subject to a steady flow of the superfluid component with a velocity of 25 cm/s along the $x$ axis. The figure shows several snapshots leading up to the annihilation (time given by the color), with 25 ns between subsequent snapshots. The entire evolution from the initial condition to the final annihilation takes about 40 $\mu$s (the entire lead-up to the reconnection is not shown).
@@
Figure 12: Annihilation of vortex half-ring with 500 nm diameter attached to the bottom wall with orientation parallel with the imposed superflow oriented along the positive x-axis direction. The initial condition was a half-circle lying in the $x=0$ plane.
@@
Figure 13: As in Fig. 12 with the ring orientation anti-parallel with the imposed $\mathbf{v}_{s}$. The vortex propagates in the negative $x$ direction because its self-induced velocity exceeds the imposed superfluid velocity.